OpenStudy (clamin):

HELP!! 5k^2+39k+28 factor

OpenStudy (anonymous):

(5k )( k )

OpenStudy (anonymous):

now you are looking for 2 number if you multiply them and when you add them gives the one in the mid

OpenStudy (anonymous):

4*7=28 2*14=28

OpenStudy (anonymous):

(5k +4 )(k +7 )

OpenStudy (anonymous):

5k*k=5k^2 5k*7=35k 4*k=4k 4*7=28 adding 5k^2+39k+28